# Moonshot: Optimizing Chain-Based Rotating Leader BFT via Optimistic Proposals

### v1

@@ Appendix B Pipelined Moonshot Security Analysis
For our induction hypothesis, we assume that the lemma holds up to view $v^{\prime}-1$. That is, we assume that every $\mathcal{C}_{v^{*}}(B_{k^{*}})$ with $v\leq v^{*}<v^{\prime}$ extends $B_{k}$. We use this assumption to prove that it also holds for $v^{\prime}$. We first observe that the existence of $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ implies that a set $H_{1}$ of at least $f+1$ honest nodes voted for $B_{k^{\prime}}$ in view $v^{\prime}$. If any of these nodes voted using the optimistic or normal vote rules then they must have received $\mathcal{C}_{v^{\prime}-1}(B_{k^{\prime}-1})$ and $B_{k^{\prime}}$ must extend $B_{k^{\prime}-1}$. Therefore, since by the induction hypothesis $B_{k^{\prime}-1}$ extends $B_{k}$, $B_{k^{\prime}}$ also extends $B_{k}$. Alternatively, if no honest node used either of these rules to vote for $B_{k^{\prime}}$ then all members of $H_{1}$ must have used the fallback vote rule to vote for $B_{k^{\prime}}$. Therefore, they must have received $\langle{\sf fb\text{-}propose},B_{k^{\prime}},\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}}),\mathcal{TC}_{v^{\prime}-1},v^{\prime}\rangle$ such that $\mathcal{C}_{v^{\prime\prime}}(B_{h})$ was the highest ranked block certificate in $\mathcal{TC}_{v^{\prime}-1}$ and $B_{k^{\prime}}$ directly extended $B_{k^{\prime\prime}}$, before multicasting a $\mathcal{T}$ message for $v^{\prime}$ or any higher view. Moreover, by 8, $v^{\prime\prime}<v^{\prime}$. We now show that $v^{\prime\prime}\geq v$.
Recall that we know from the commit rule that $\mathcal{C}_{v+1}(B_{k+1})$ exists and that $B_{k+1}$ extends $B_{k}$. Therefore, a set $H_{2}$ of at least $f+1$ honest nodes must have voted for $B_{k+1}$ in view $v+1$. By the vote rules and the lock rule, these nodes must have received $\mathcal{C}_{v}(B_{k})$ and must not have sent $\mathcal{T}_{v^{*}}$ with $v^{*}\geq v+1$ before doing so. Thus, by 1, every $\mathcal{T}_{v^{*}}$ message sent by $H_{2}$ will necessarily contain $\mathcal{C}_{v}(B_{k})$ or higher. Therefore, because $v^{\prime}-1\geq v+1$ and since by 2 $H_{1}$ and $H_{2}$ must have at least one node in common, the highest ranked block certificate of $\mathcal{TC}_{v^{\prime}-1}$ must have a rank of at least as great as $\mathcal{C}_{v}(B_{k})$. Thus, $v^{\prime\prime}\geq v$. Therefore, since $v\leq v^{\prime\prime}<v^{\prime}$, by the induction hypothesis, $B_{k^{\prime\prime}}$ extends $B_{k}$, so $B_{k^{\prime}}$ also extends $B_{k}$. ∎
@@
The view advancement rule requires an honest node to receive either $\mathcal{C}_{v-1}$ or $\mathcal{TC}_{v-1}$ in order to enter $v$. Therefore, at least $f+1$ honest nodes must multicast the corresponding messages. In the case of $\mathcal{C}_{v-1}$, the vote rules require these nodes to be in $v-1$ when they do so. In the case of $\mathcal{TC}_{v-1}$, at least one honest node must have its view timer expire whilst in $v-1$ before any honest node can multicast $\mathcal{T}_{v-1}$. Thus, in either case, an honest node can only enter $v$ if at least one honest node has already entered $v-1$. ∎
@@
If any honest node enters view $v^{\prime}$ such that $v^{\prime}\geq v$ via $\mathcal{C}_{v^{\prime}-1}$ before $t+\Delta$ then, by the view advancement rules, it will multicast $\mathcal{C}_{v^{\prime}-1}$ and thus all honest nodes will enter $v^{\prime}$ or higher before $t+2\Delta$. Suppose, then, that no honest node enters $v^{\prime}$ via $\mathcal{C}_{v^{\prime}-1}$ before $t+\Delta$. Therefore, $P_{i}$ must enter $v$ via $\mathcal{TC}_{v-1}$. Hence, at least $f+1$ honest nodes must multicast $\mathcal{T}_{v-1}$ before $t$ and thus all will receive these messages before $t+\Delta$. Furthermore, since $t$ is defined as the time that the first honest node enters $v$, by Corollary 1, $\mathcal{TC}_{v^{\prime}}$ cannot exist before $t+3\Delta$ and thus all honest nodes must be in view $v$ or lower when they receive the aforementioned $\mathcal{T}_{v-1}$ messages. Hence, by the timeout rule, all honest nodes in $v-1$ or lower that have not already multicasted $\mathcal{T}_{v-1}$ will do so before $t+\Delta$. Moreover, every honest node that enters $v$ before this time must do so via $\mathcal{TC}_{v-1}$ and thus, by the timeout rule, will also multicast $\mathcal{T}_{v-1}$ before $t+\Delta$. Consequently, all honest nodes will multicast $\mathcal{T}_{v-1}$ before this time, so they will all be able to construct $\mathcal{TC}_{v-1}$ before $t+2\Delta$. Thus, by the round transition rules, every honest node will enter $v$ or higher before $t+2\Delta$. ∎
@@
By Lemma 7, only one block can become certified for a given view. Thus, if $\mathcal{C}_{v}(B_{k})$ exists then any node that receives a view $v$ block certificate must receive $\mathcal{C}_{v}(B_{k})$. Additionally, by Lemma 9 and 11, all honest nodes enter $v$ or higher and receive a proposal from $L_{v}$ before $t+2\Delta$. Moreover, since $t$ is defined as the time that the first honest node enters $v$, no honest node will multicast $\mathcal{T}_{v}$ before $t+3\Delta$. Therefore, if any honest node enters a view greater than $v$ before $t+2\Delta$ then, by 9, at least one honest node must have already entered $v+1$ via $\mathcal{C}_{v}(B_{k})$. By the view advancement rule, this node would have multicasted $\mathcal{C}_{v}(B_{k})$, so all nodes will receive this certificate before $t+3\Delta$, completing the proof. Alternatively, if no honest node receives $\mathcal{C}_{v}(B_{k})$ before $t+2\Delta$, then all honest nodes will enter $v$ before $t+2\Delta$. Moreover, since $L_{v}$ is honest, it will ensure that its proposal is well-formed: i.e., if it is a normal proposal then the proposed block will extend the block certified by the included block certificate; if it is a fallback proposal then the proposed block will extend the block certified by the block certificate with the highest rank in the included timeout certificate. Additionally, since $L_{v}$ is honest, it will create only one normal proposal or one fallback proposal. Moreover, if it creates a normal proposal then any equivocal optimistic proposal that it may have created will necessarily have a different parent than the normal proposal because honest leaders propose fixed block payloads for a given view, per the definition of a block given in Section II. Consequently, by Lemma 7, the parent of the equivocal optimistic block proposal cannot be certified, so, by the optimistic vote rule, no honest node will be able to vote for this proposal. Finally, since all honest nodes will receive $L_{v}$’s proposal before $t+2\Delta$, they cannot have ${\sf timeout\_view}_{i}\geq v$ by this time. Thus, by the vote rules, they will all vote for the included block, so all honest nodes will be able to construct $\mathcal{C}_{v}(B_{k})$ before $t+3\Delta$. ∎
@@
For our induction hypothesis, we assume that the theorem holds up to view $v^{\prime}-1$. That is, we assume that every $\mathcal{C}_{v^{*}}(B_{k^{*}})$ with $v\leq v^{*}<v^{\prime}$ extends $B_{k}$. We use this assumption to prove that it also holds for $v^{\prime}$. We first observe that the existence of $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ implies that a set $H_{1}$ of at least $f+1$ honest nodes voted for $B_{k^{\prime}}$ in view $v^{\prime}$. If any of these nodes voted using the optimistic or normal vote rules then they must have received $\mathcal{C}_{v^{\prime}-1}(B_{k^{\prime}-1})$ and $B_{k^{\prime}}$ must extend $B_{k^{\prime}-1}$. Therefore, since by the induction hypothesis $B_{k^{\prime}-1}$ extends $B_{k}$, $B_{k^{\prime}}$ also extends $B_{k}$. Alternatively, if no honest node used either of these rules to vote for $B_{k^{\prime}}$ then all members of $H_{1}$ must have used the fallback vote rule to vote for $B_{k^{\prime}}$. Therefore, they must have received $\langle{\sf fb\text{-}propose},B_{k^{\prime}},\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}}),\mathcal{TC}_{v^{\prime}-1},v^{\prime}\rangle$ such that $\mathcal{C}_{v^{\prime\prime}}(B_{h})$ was the highest ranked block certificate in $\mathcal{TC}_{v^{\prime}-1}$ and $B_{k^{\prime}}$ directly extended $B_{k^{\prime\prime}}$, before multicasting a $\mathcal{T}$ message for $v^{\prime}$ or any higher view. Moreover, by 8, $v^{\prime\prime}<v^{\prime}$. We now show that $v^{\prime\prime}\geq v$.
By the fallback vote rule, $H_{1}$ will only vote for a fallback proposal for $v^{\prime}$ if it contains a valid $\mathcal{TC}_{v^{\prime}-1}$. Hence, a set $H_{2}$ of at least $f+1$ honest nodes must multicast $\mathcal{T}_{v^{\prime}-1}$. Moreover, by Lemma 13, a set $H_{3}$ of at least $f+1$ honest nodes must lock $\mathcal{C}_{v}(B_{k})$ while entering $v+1$ and must do so without having multicasted $\mathcal{T}_{v^{\prime}}$ for $v^{\prime}\geq v$. By 2, $H_{2}$ and $H_{3}$ must have at least one node, say $P_{i}$, in common. Therefore, since $P_{i}$ must lock $\mathcal{C}_{v}(B_{k})$ before sending $\mathcal{T}_{v^{\prime}-1}$, by the lock rule and the block certificate ranking rule, every valid $\mathcal{TC}_{v^{\prime}-1}$ must contain a block certificate with a rank of at least $v$. Thus, $\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}})\geq\mathcal{C}_{v}(B_{k})$, so $v^{\prime\prime}\geq v$. Therefore, since $v\leq v^{\prime\prime}<v^{\prime}$, by the induction hypothesis, $B_{k^{\prime}}$ extends $B_{k}$. ∎
@@ Appendix C Commit Moonshot Security Analysis
We provide an updated proof for Lemma 8 below, which, when invoked in place of the original lemma in Theorem 6, makes Theorem 6 a sufficient proof of the safety of Commit Moonshot. The liveness and reorg resilience proofs for Commit Moonshot remain identical to those given in Theorem 7 and Theorem 8.
Lemma 15 (Unique Extensibility).
@@
By the direct commit rule, $P_{i}$ must have received $12f+$ 1 $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$, at least $f+1$ of which must have been sent by a set $H$ of distinct honest nodes. Hence, by the pre-commit rule, $H$ must have received $\mathcal{C}_{v}(B_{k})$ whilst having ${\sf timeout\_view}_{i}<v$. Consequently, by the view advancement rule, $H$ must have entered $v+1$ before sending $\mathsf{timeout}_{v}$ and thus, by the timeout rule, will never send $\mathsf{timeout}_{v}$. Therefore, since $\mathcal{TC}_{v}$ cannot exist, $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ must be either $\mathcal{C}^{o}_{v+1}(B_{k^{\prime}})$ or $\mathcal{C}^{n}_{v+1}(B_{k^{\prime}})$. In either case, by the respective vote rules, $B_{k^{\prime}}$ extends $B_{k}$.
@@
For our induction hypothesis, we assume that the lemma holds up to view $v^{\prime}-1$. That is, we assume that every $\mathcal{C}_{v^{*}}(B_{k^{*}})$ with $v\leq v^{*}<v^{\prime}$ extends $B_{k}$. We use this assumption to prove that it also holds for $v^{\prime}$. We first observe that the existence of $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ implies that a set $H_{1}$ of at least $f+1$ honest nodes voted for $B_{k^{\prime}}$ in view $v^{\prime}$. If any of these nodes voted using the optimistic or normal vote rules then they must have observed $\mathcal{C}_{v^{\prime}-1}(B_{k^{\prime}-1})$ and $B_{k^{\prime}}$ must extend $B_{k^{\prime}-1}$. Therefore, since by the induction hypothesis $B_{k^{\prime}-1}$ extends $B_{k}$, $B_{k^{\prime}}$ also extends $B_{k}$. Alternatively, if no honest node used either of these rules to vote for $B_{k^{\prime}}$ then all members of $H_{1}$ must have used the fallback vote rule to vote for $B_{k^{\prime}}$. Therefore, they must have received $\langle{\sf fb\text{-}propose},B_{k^{\prime}},\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}}),\mathcal{TC}_{v^{\prime}-1},v^{\prime}\rangle$ such that $\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}})$ was the highest ranked block certificate in $\mathcal{TC}_{v^{\prime}-1}$ and $B_{k^{\prime}}$ directly extended $B_{k^{\prime\prime}}$, before multicasting $\mathcal{T}$ for $v^{\prime}$ or any higher view. Moreover, by 8, $v^{\prime\prime}<v^{\prime}$. We now show that $v^{\prime\prime}\geq v$.

### v2

@@ Appendix B Pipelined Moonshot Security Analysis
For our induction hypothesis, we assume that the lemma holds up to view $v^{\prime}-1$. That is, we assume that every $\mathcal{C}_{v^{*}}(B_{k^{*}})$ with $v\leq v^{*}<v^{\prime}$ extends $B_{k}$. We use this assumption to prove that it also holds for $v^{\prime}$. We first observe that the existence of $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ implies that a set $H_{1}$ of at least $f+1$ honest nodes voted for $B_{k^{\prime}}$ in view $v^{\prime}$. If any of these nodes voted using the optimistic or normal vote rules then they must have received $\mathcal{C}_{v^{\prime}-1}(B_{k^{\prime}-1})$ and $B_{k^{\prime}}$ must extend $B_{k^{\prime}-1}$. Therefore, since by the induction hypothesis $B_{k^{\prime}-1}$ extends $B_{k}$, $B_{k^{\prime}}$ also extends $B_{k}$. Alternatively, if no honest node used either of these rules to vote for $B_{k^{\prime}}$ then all members of $H_{1}$ must have used the fallback vote rule to vote for $B_{k^{\prime}}$. Therefore, they must have received $\langle{\sf fb\text{-}propose},B_{k^{\prime}},\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}}),\mathcal{TC}_{v^{\prime}-1},v^{\prime}\rangle$ such that $\mathcal{C}_{v^{\prime\prime}}(B_{h})$ had a rank greater than or equal to that of the highest ranked block certificate in $\mathcal{TC}_{v^{\prime}-1}$ and $B_{k^{\prime}}$ directly extended $B_{k^{\prime\prime}}$, before multicasting a $\mathcal{T}$ message for $v^{\prime}$ or any higher view. Moreover, by 8, $v^{\prime\prime}<v^{\prime}$. We now show that $v^{\prime\prime}\geq v$.
Recall that we know from the commit rule that $\mathcal{C}_{v+1}(B_{k+1})$ exists and that $B_{k+1}$ extends $B_{k}$. Therefore, a set $H_{2}$ of at least $f+1$ honest nodes must have voted for $B_{k+1}$ in view $v+1$. By the vote rules and the lock rule, these nodes must have received $\mathcal{C}_{v}(B_{k})$ and must not have sent $\mathcal{T}_{v^{*}}$ with $v^{*}\geq v+1$ before doing so. Thus, by 1, every $\mathcal{T}_{v^{*}}$ message sent by $H_{2}$ will necessarily contain $\mathcal{C}_{v}(B_{k})$ or higher. Therefore, because $v^{\prime}-1\geq v+1$ and since by 2 $H_{1}$ and $H_{2}$ must have at least one node in common, the highest ranked block certificate of $\mathcal{TC}_{v^{\prime}-1}$ must have a rank of at least as great as $\mathcal{C}_{v}(B_{k})$. Therefore, by extension and the definition of the rank of a block certificate given in Section II, $v^{\prime\prime}\geq v$. Therefore, since $v\leq v^{\prime\prime}<v^{\prime}$, by the induction hypothesis, $B_{k^{\prime\prime}}$ extends $B_{k}$, so $B_{k^{\prime}}$ also extends $B_{k}$. ∎
@@
The view advancement rule requires an honest node to receive either $\mathcal{C}_{v-1}$ or $\mathcal{TC}_{v-1}$ in order to enter $v$. Therefore, at least $f+1$ honest nodes must multicast the corresponding constituent messages. In the case of $\mathcal{C}_{v-1}$, the vote rules require these nodes to be in $v-1$ when they do so. In the case of $\mathcal{TC}_{v-1}$, at least one honest node must have its view timer expire whilst in $v-1$ before any honest node can multicast $\mathcal{T}_{v-1}$. Thus, in either case, an honest node can only enter $v$ if at least one honest node has already entered $v-1$. ∎
@@
If any honest node enters view $v^{\prime}$ such that $v^{\prime}\geq v$ via $\mathcal{C}_{v^{\prime}-1}$ before $t+\Delta$ then, by the view advancement rules, it will multicast $\mathcal{C}_{v^{\prime}-1}$ and thus all honest nodes will enter $v^{\prime}$ or higher before $t+2\Delta$. Suppose, then, that no honest node enters $v^{\prime}$ via $\mathcal{C}_{v^{\prime}-1}$ before $t+\Delta$. Therefore, $P_{i}$ must enter $v$ via $\mathcal{TC}_{v-1}$. Hence, at least $f+1$ honest nodes must multicast $\mathcal{T}_{v-1}$ before $t$ and thus all will receive these messages before $t+\Delta$. Furthermore, since $t$ is defined as the time that the first honest node enters $v$, by Corollary 1, $\mathcal{TC}_{v^{\prime}}$ cannot exist before $t+3\Delta$ and thus all honest nodes must be in view $v$ or lower when they receive the aforementioned $\mathcal{T}_{v-1}$ messages. Hence, by the timeout rule, all honest nodes in $v-1$ or lower that have not already multicasted $\mathcal{T}_{v-1}$ will do so before $t+\Delta$. Moreover, every honest node that enters $v$ before this time must do so via $\mathcal{TC}_{v-1}$ and thus, by the timeout rule, will also multicast $\mathcal{T}_{v-1}$ before $t+\Delta$. Consequently, all honest nodes will multicast $\mathcal{T}_{v-1}$ before this time, so they will all be able to construct $\mathcal{TC}_{v-1}$ before $t+2\Delta$. Thus, by the view transition rules, every honest node will enter $v$ or higher before $t+2\Delta$. ∎
@@
By Lemma 7, only one block can become certified for a given view. Thus, if $\mathcal{C}_{v}(B_{k})$ exists then any node that receives a view $v$ block certificate must receive $\mathcal{C}_{v}(B_{k})$. Additionally, by Lemma 9 and 11, all honest nodes enter $v$ or higher and receive a proposal from $L_{v}$ before $t+2\Delta$. Moreover, since $t$ is defined as the time that the first honest node enters $v$, no honest node will multicast $\mathcal{T}_{v}$ before $t+3\Delta$. Therefore, if any honest node enters a view greater than $v$ before $t+2\Delta$ then, by 9, at least one honest node must have already entered $v+1$ via $\mathcal{C}_{v}(B_{k})$. By the view advancement rule, this node would have multicasted $\mathcal{C}_{v}(B_{k})$, so all nodes will receive this certificate before $t+3\Delta$, completing the proof. Alternatively, if no honest node receives $\mathcal{C}_{v}(B_{k})$ before $t+2\Delta$, then all honest nodes will enter $v$ before $t+2\Delta$. Moreover, since $L_{v}$ is honest, it will ensure that its proposal is well-formed: i.e., if it is a normal proposal then the proposed block will extend the block certified by the included block certificate. Otherwise, if it is a fallback proposal then the proposed block will extend the block certified by $L_{v}$’s ${\sf lock}_{i}$, which, by the lock rule, will have a rank at least as great as that of the block certificate with the highest rank in the included timeout certificate. Additionally, since $L_{v}$ is honest, it will create only one normal proposal or one fallback proposal. Moreover, if it creates a normal proposal then any equivocal optimistic proposal that it may have created will necessarily have a different parent than the normal proposal because honest leaders propose fixed block payloads for a given view, per the definition of a block given in Section II. Consequently, by Lemma 7, the parent of the equivocal optimistic block proposal cannot be certified, so, by the optimistic vote rule, no honest node will be able to vote for this proposal. Finally, since all honest nodes will receive $L_{v}$’s proposal before $t+2\Delta$, they cannot have ${\sf timeout\_view}_{i}\geq v$ by this time. Thus, by the vote rules, they will all vote for the included block, so all honest nodes will be able to construct $\mathcal{C}_{v}(B_{k})$ before $t+3\Delta$. ∎
@@
For our induction hypothesis, we assume that the theorem holds up to view $v^{\prime}-1$. That is, we assume that every $\mathcal{C}_{v^{*}}(B_{k^{*}})$ with $v\leq v^{*}<v^{\prime}$ extends $B_{k}$. We use this assumption to prove that it also holds for $v^{\prime}$. We first observe that the existence of $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ implies that a set $H_{1}$ of at least $f+1$ honest nodes voted for $B_{k^{\prime}}$ in view $v^{\prime}$. If any of these nodes voted using the optimistic or normal vote rules then they must have received $\mathcal{C}_{v^{\prime}-1}(B_{k^{\prime}-1})$ and $B_{k^{\prime}}$ must extend $B_{k^{\prime}-1}$. Therefore, since by the induction hypothesis $B_{k^{\prime}-1}$ extends $B_{k}$, $B_{k^{\prime}}$ also extends $B_{k}$. Alternatively, if no honest node used either of these rules to vote for $B_{k^{\prime}}$ then all members of $H_{1}$ must have used the fallback vote rule to vote for $B_{k^{\prime}}$. Therefore, they must have received $\langle{\sf fb\text{-}propose},B_{k^{\prime}},\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}}),\mathcal{TC}_{v^{\prime}-1},v^{\prime}\rangle$ such that $\mathcal{C}_{v^{\prime\prime}}(B_{h})$ had a rank greater than or equal to that of the highest ranked block certificate in $\mathcal{TC}_{v^{\prime}-1}$ and $B_{k^{\prime}}$ directly extended $B_{k^{\prime\prime}}$, before multicasting a $\mathcal{T}$ message for $v^{\prime}$ or any higher view. Moreover, by 8, $v^{\prime\prime}<v^{\prime}$. We now show that $v^{\prime\prime}\geq v$.
By the fallback vote rule, $H_{1}$ will not vote for a fallback proposal for $v^{\prime}$ unless it contains a valid $\mathcal{TC}_{v^{\prime}-1}$. Hence, a set $H_{2}$ of at least $f+1$ honest nodes must multicast $\mathcal{T}_{v^{\prime}-1}$. Moreover, by Lemma 13, a set $H_{3}$ of at least $f+1$ honest nodes must lock $\mathcal{C}_{v}(B_{k})$ while entering $v+1$ and must do so without having multicasted $\mathcal{T}_{v^{\prime}}$ for $v^{\prime}\geq v$. By 2, $H_{2}$ and $H_{3}$ must have at least one node, say $P_{i}$, in common. Therefore, since $P_{i}$ must lock $\mathcal{C}_{v}(B_{k})$ before sending $\mathcal{T}_{v^{\prime}-1}$, by the lock rule and the block certificate ranking rule, every valid $\mathcal{TC}_{v^{\prime}-1}$ must contain a block certificate with a rank of at least $v$. Thus, $\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}})\geq\mathcal{C}_{v}(B_{k})$, so $v^{\prime\prime}\geq v$. Therefore, since $v\leq v^{\prime\prime}<v^{\prime}$, by the induction hypothesis, $B_{k^{\prime}}$ extends $B_{k}$. ∎
@@ Appendix C Commit Moonshot Security Analysis
As previously observed, because Commit Moonshot retains the vote and commit rules of Pipelined Moonshot, the same liveness argument that can be made for the latter also applies to the former. The same is also true for the reorg resilience of the protocol. However, Commit Moonshot’s new commit path requires additional justification. We prove the safety of the additional rules below, and observe that Theorem 4 holds for Commit Moonshot when Lemma 15 is invoked along with Lemma 8. We subsequently prove that Commit Moonshot requires only a single honest leader to commit a new block after GST.
Claim 12.
If $12f+$ 1 distinct nodes send $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ then $\mathcal{TC}_{v}$ cannot exist.
Proof.
Let $H$ denote the set of $f+1$ honest nodes that sent $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$. By the pre-commit rules, any member of $H$ that sent this message must have received $\mathcal{C}_{v}(B_{k})$ whilst having ${\sf timeout\_view}_{*}<v$. Therefore, by the view advancement rule, all members of $H$ must have entered $v+1$ before sending $\mathsf{timeout}_{v}$ and thus, by the timeout rule, will never send $\mathsf{timeout}_{v}$. Therefore, since this implies that at most $2f$ distinct $\mathsf{timeout}_{v}$ messages will ever exist, $\mathcal{TC}_{v}$ cannot exist. ∎
Claim 13.
If $12f+$ 1 distinct nodes send $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ then every $\mathcal{TC}_{v^{\prime}}$ for view $v^{\prime}>v$ must contain a block certificate for $v$ or higher.
Proof.
Let $H$ denote the set of $f+1$ honest nodes that sent $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$. By the lock rule, the members of $H$ that voted to commit $B_{k}$ via the direct pre-commit rule must have locked $\mathcal{C}_{v}(B_{k})$ upon receiving it. Comparatively, those that voted to commit $B_{k}$ via the indirect pre-commit rule must have previously sent $\langle\mathsf{commit},H(B_{l}),v^{\prime\prime}\rangle_{*}$ for some descendant $B_{l}$ of $B_{k}$ for some $v^{\prime\prime}>v$. Thus, by the pre-commit and lock rules, these nodes must have received and locked $\mathcal{C}_{v^{\prime\prime}}(B_{l})$ or some higher ranked block certificate before sending $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$. In either case, the members of $H$ must have locked a block certificate for $v$ or higher whilst having ${\sf timeout\_view}_{*}<v$, so all timeout messages sent by these nodes for views greater than $v$ will necessarily contain $\mathcal{C}_{v}(B_{k})$ or higher. Thus, since every $\mathcal{TC}_{v^{\prime}}$ for view $v^{\prime}>v$ must contain a timeout message from at least one member of $H$, every such certificate must contain a block certificate for $v$ or higher. ∎
Lemma 15 (Unique Extensibility Continued).
If an honest node, say $P_{i}$, directly commits a block $B_{k}$ via the alternative direct commit rule that was certified for view $v$ and $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ exists such that $v^{\prime}\geq v$, then $B_{k^{\prime}}$ extends $B_{k}$.
@@
By 12, $\mathcal{TC}_{v}$ cannot exist. Therefore, $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ must be either $\mathcal{C}^{o}_{v+1}(B_{k^{\prime}})$ or $\mathcal{C}^{n}_{v+1}(B_{k^{\prime}})$. In either case, by the respective vote rules, $B_{k^{\prime}}$ extends $B_{k}$.
@@
For our induction hypothesis, we assume that the lemma holds up to view $v^{\prime}-1$. That is, we assume that every $\mathcal{C}_{v^{*}}(B_{k^{*}})$ with $v\leq v^{*}<v^{\prime}$ extends $B_{k}$. We use this assumption to prove that it also holds for $v^{\prime}$. We first observe that the existence of $\mathcal{C}_{v^{\prime}}(B_{k^{\prime}})$ implies that a set $H_{1}$ of at least $f+1$ honest nodes voted for $B_{k^{\prime}}$ in view $v^{\prime}$. If any of these nodes voted using the optimistic or normal vote rules then they must have observed $\mathcal{C}_{v^{\prime}-1}(B_{k^{\prime}-1})$ and $B_{k^{\prime}}$ must extend $B_{k^{\prime}-1}$. Therefore, since by the induction hypothesis $B_{k^{\prime}-1}$ extends $B_{k}$, $B_{k^{\prime}}$ also extends $B_{k}$. Alternatively, if no honest node used either of these rules to vote for $B_{k^{\prime}}$ then all members of $H_{1}$ must have used the fallback vote rule to vote for $B_{k^{\prime}}$. Therefore, they must have received $\langle{\sf fb\text{-}propose},B_{k^{\prime}},\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}}),\mathcal{TC}_{v^{\prime}-1},v^{\prime}\rangle$ such that $\mathcal{C}_{v^{\prime\prime}}(B_{k^{\prime\prime}})$ had a rank at least as great as that of the highest ranked block certificate in $\mathcal{TC}_{v^{\prime}-1}$ and $B_{k^{\prime}}$ directly extended $B_{k^{\prime\prime}}$, before multicasting $\mathcal{T}$ for $v^{\prime}$ or any higher view. Moreover, by 8, $v^{\prime\prime}<v^{\prime}$. We now show that $v^{\prime\prime}\geq v$.